\begin{thebibliography}{10}

\bibitem{assiotis2022infinite}
Theodoros Assiotis.
\newblock Infinite p-adic random matrices and ergodic decomposition of p-adic {H}ua measures.
\newblock {\em Transactions of the American Mathematical Society}, 375(3):1745--1766, 2022.

\bibitem{band1}
Zhigang Bao and L{\'a}szl{\'o} Erd{\H{o}}s.
\newblock Delocalization for a class of random block band matrices.
\newblock {\em Probability Theory and Related Fields}, 167:673--776, 2017.

\bibitem{bourgade2018random}
Paul Bourgade.
\newblock Random band matrices.
\newblock In {\em Proceedings of the International Congress of Mathematicians: Rio de Janeiro 2018}, pages 2759--2783. World Scientific, 2018.

\bibitem{bb2}
Paul Bourgade, Fan Yang, H-T Yau, and Jun Yin.
\newblock Random band matrices in the delocalized phase, {II}: Generalized resolvent estimates.
\newblock {\em Journal of Statistical Physics}, 174:1189--1221, 2019.

\bibitem{bb1}
Paul Bourgade, Horng~Tzer Yau, and Jun Yin.
\newblock Random band matrices in the delocalized phase {I}: Quantum unique ergodicity and universality.
\newblock {\em Communications on Pure and Applied Mathematics}, 73(7):1526--1596, 2020.

\bibitem{brodie2022density}
Benjamin Brodie and Peter~D Hislop.
\newblock The density of states and local eigenvalue statistics for random band matrices of fixed width.
\newblock {\em Journal of Spectral Theory}, 12(2):417--457, 2022.

\bibitem{bufetov2017ergodic}
Alexander~I Bufetov and Yanqi Qiu.
\newblock Ergodic measures on spaces of infinite matrices over non-{A}rchimedean locally compact fields.
\newblock {\em Compositio Mathematica}, 153(12):2482--2533, 2017.

\bibitem{chen2022random}
Nixia Chen and Charles~K Smart.
\newblock Random band matrix localization by scalar fluctuations.
\newblock {\em arXiv preprint arXiv:2206.06439}, 2022.

\bibitem{recent4}
Gilyoung Cheong and Nathan Kaplan.
\newblock Generalizations of results of {F}riedman and {W}ashington on cokernels of random p-adic matrices.
\newblock {\em Journal of Algebra}, 604:636--663, 2022.

\bibitem{recent6}
Gilyoung Cheong, Yunqi Liang, and Michael Strand.
\newblock The distribution of the cokernel of a polynomial push-forward of a random $\mathbb{Z}_p$-matrix with a fixed residue class modulo $ p$.
\newblock {\em arXiv preprint arXiv:2209.03626}, 2022.

\bibitem{recent7}
Gilyoung Cheong and Myungjun Yu.
\newblock The cokernel of a polynomial of a random integral matrix.
\newblock {\em arXiv preprint arXiv:2303.09125}, 2023.

\bibitem{cipolloni2024dynamical}
Giorgio Cipolloni, Ron Peled, Jeffrey Schenker, and Jacob Shapiro.
\newblock Dynamical localization for random band matrices up to $w\ll n^{1/4}$.
\newblock {\em Communications in Mathematical Physics}, 405(3):82, 2024.

\bibitem{clancy2015cohen}
Julien Clancy, Nathan Kaplan, Timothy Leake, Sam Payne, and Melanie~Matchett Wood.
\newblock On a {C}ohen--{L}enstra heuristic for {J}acobians of random graphs.
\newblock {\em Journal of Algebraic Combinatorics}, 42:701--723, 2015.

\bibitem{clancy2015note}
Julien Clancy, Timothy Leake, and Sam Payne.
\newblock A note on {J}acobians, {T}utte polynomials, and two-variable zeta functions of graphs.
\newblock {\em Experimental Mathematics}, 24(1):1--7, 2015.

\bibitem{cohen2006heuristics}
Henri Cohen and Hendrik~W Lenstra~Jr.
\newblock Heuristics on class groups of number fields.
\newblock In {\em Number Theory Noordwijkerhout 1983: Proceedings of the Journ{\'e}es Arithm{\'e}tiques held at Noordwijkerhout, The Netherlands July 11--15, 1983}, pages 33--62. Springer, 2006.

\bibitem{ellenberg2016homological}
Jordan~S Ellenberg, Akshay Venkatesh, and Craig Westerland.
\newblock Homological stability for {H}urwitz spaces and the {C}ohen-{L}enstra conjecture over function fields.
\newblock {\em Annals of Mathematics}, pages 729--786, 2016.

\bibitem{band2}
L{\'a}szl{\'o} Erd{\H{o}}s and Antti Knowles.
\newblock Quantum diffusion and delocalization for band matrices with general distribution.
\newblock In {\em Annales Henri Poincar{\'e}}, volume~12, pages 1227--1319. Springer, 2011.

\bibitem{band3}
L{\'a}szl{\'o} Erd{\H{o}}s and Antti Knowles.
\newblock The {A}ltshuler--{S}hklovskii formulas for random band matrices {I}: the unimodular case.
\newblock {\em Communications in Mathematical Physics}, 333(3):1365--1416, 2015.

\bibitem{band4}
L{\'a}szl{\'o} Erd{\H{o}}s and Antti Knowles.
\newblock The {A}ltshuler--{S}hklovskii formulas for random band matrices {II}: the general case.
\newblock In {\em Annales Henri Poincar{\'e}}, volume~16, pages 709--799. Springer, 2015.

\bibitem{band5}
L{\'a}szl{\'o} Erd{\H{o}}s, Antti Knowles, Horng-Tzer Yau, and Jun Yin.
\newblock Delocalization and diffusion profile for random band matrices.
\newblock {\em Communications in Mathematical Physics}, 323:367--416, 2013.

\bibitem{friedman1987distribution}
Eduardo Friedman and Lawrence~C Washington.
\newblock On the distribution of divisor class groups of curves over a finite field.
\newblock {\em Th{\'e}orie des nombres (Quebec, PQ, 1987)}, pages 227--239, 1987.

\bibitem{recent9}
Elia Gorokhovsky.
\newblock Time-inhomogeneous random walks on finite groups and cokernels of random integer block matrices.
\newblock {\em arXiv preprint arXiv:2405.11435}, 2024.

\bibitem{hislop2022local}
Peter~D Hislop and M~Krishna.
\newblock On the local eigenvalue statistics for random band matrices in the localization regime.
\newblock {\em Journal of Statistical Physics}, 187(3):26, 2022.

\bibitem{recent8}
Eliot Hodges.
\newblock The distribution of sandpile groups of random graphs with their pairings.
\newblock {\em arXiv preprint arXiv:2311.07078}, 2023.

\bibitem{recent3}
Jiwan Jung and Jungin Lee.
\newblock Joint distribution of the cokernels of random $ p $-adic matrices {II}.
\newblock {\em arXiv preprint arXiv:2304.03583}, 2023.

\bibitem{recent2}
Jungin Lee.
\newblock Joint distribution of the cokernels of random p-adic matrices.
\newblock In {\em Forum Mathematicum}, number~0. De Gruyter, 2023.

\bibitem{meszaros2020distribution}
Andr{\'a}s M{\'e}sz{\'a}ros.
\newblock The distribution of sandpile groups of random regular graphs.
\newblock {\em Transactions of the American Mathematical Society}, 373(9):6529--6594, 2020.

\bibitem{meszaros2023cohen}
Andr{\'a}s M{\'e}sz{\'a}ros.
\newblock {C}ohen-{L}enstra distribution for sparse matrices with determinantal biasing.
\newblock {\em arXiv preprint arXiv:2307.04741}, 2023.

\bibitem{meszaros20242}
Andr{\'a}s M{\'e}sz{\'a}ros.
\newblock The $2 $-torsion of determinantal hypertrees is not {C}ohen-{L}enstra.
\newblock {\em arXiv preprint arXiv:2404.02308}, 2024.

\bibitem{neretin2013hua}
Yu~A Neretin.
\newblock Hua measures on the space of p-adic matrices and inverse limits of {G}rassmannians.
\newblock {\em Izvestiya. Mathematics}, 77(5):941--953, 2013.

\bibitem{recent1}
Hoi~H Nguyen and Roger Van~Peski.
\newblock Universality for cokernels of random matrix products.
\newblock {\em arXiv preprint arXiv:2209.14957}, 2022.

\bibitem{recent10}
Hoi~H Nguyen and Melanie~Matchett Wood.
\newblock Local and global universality of random matrix cokernels.
\newblock {\em arXiv preprint arXiv:2210.08526}, 2022.

\bibitem{nguyen2022random}
Hoi~H Nguyen and Melanie~Matchett Wood.
\newblock Random integral matrices: universality of surjectivity and the cokernel.
\newblock {\em Inventiones mathematicae}, 228(1):1--76, 2022.

\bibitem{peled2019wegner}
Ron Peled, Jeffrey Schenker, Mira Shamis, and Sasha Sodin.
\newblock On the {W}egner orbital model.
\newblock {\em International Mathematics Research Notices}, 2019(4):1030--1058, 2019.

\bibitem{sawin2022moment}
Will Sawin and Melanie~Matchett Wood.
\newblock The moment problem for random objects in a category.
\newblock {\em arXiv preprint arXiv:2210.06279}, 2022.

\bibitem{schenker2009eigenvector}
Jeffrey Schenker.
\newblock Eigenvector localization for random band matrices with power law band width.
\newblock {\em Communications in Mathematical Physics}, 290:1065--1097, 2009.

\bibitem{van2021limits}
Roger Van~Peski.
\newblock Limits and fluctuations of p-adic random matrix products.
\newblock {\em Selecta Mathematica}, 27:1--71, 2021.

\bibitem{recent5}
Roger Van~Peski.
\newblock {H}all--{L}ittlewood polynomials, boundaries, and p-adic random matrices.
\newblock {\em International Mathematics Research Notices}, 2023(13):11217--11275, 2023.

\bibitem{van2023local}
Roger Van~Peski.
\newblock Local limits in $ p $-adic random matrix theory.
\newblock {\em arXiv preprint arXiv:2310.12275}, 2023.

\bibitem{van2023p}
Roger Van~Peski.
\newblock What is a $ p $-adic {D}yson {B}rownian motion?
\newblock {\em arXiv preprint arXiv:2309.02865}, 2023.

\bibitem{van2024symmetric}
Roger Van~Peski.
\newblock Symmetric functions and the explicit moment problem for abelian groups.
\newblock {\em arXiv preprint arXiv:2402.16625}, 2024.

\bibitem{wood2017distribution}
Melanie Wood.
\newblock The distribution of sandpile groups of random graphs.
\newblock {\em Journal of the American Mathematical Society}, 30(4):915--958, 2017.

\bibitem{wood2019random}
Melanie~Matchett Wood.
\newblock Random integral matrices and the {C}ohen-{L}enstra heuristics.
\newblock {\em American Journal of Mathematics}, 141(2):383--398, 2019.

\bibitem{wood2022probability}
Melanie~Matchett Wood.
\newblock Probability theory for random groups arising in number theory.
\newblock In {\em Proceedings of the International Congress of Mathematicians}, pages 4476--4508, 2022.

\bibitem{bb3}
Fan Yang and Jun Yin.
\newblock Random band matrices in the delocalized phase, {III}: Averaging fluctuations.
\newblock {\em Probability Theory and Related Fields}, 179:451--540, 2021.

\end{thebibliography}

\documentclass[12pt,a4paper]{article}

\usepackage{authblk}
\usepackage[margin=2cm]{geometry}
\usepackage{t1enc}
\usepackage[utf8]{inputenc}
\usepackage{amsthm,amsmath,amssymb}
\usepackage{graphicx}
\usepackage{enumerate}
\usepackage{hyperref}
\usepackage{bm}
\usepackage{comment}
\usepackage{amsfonts}
\usepackage{graphicx,caption}
\usepackage{bm}
\usepackage{amsmath, amsthm, amssymb}
\usepackage{graphicx}
\usepackage{hyperref}
\usepackage{relsize}
\usepackage{blkarray}
\usepackage{algpseudocode}
\newcommand{\lC}{L}

\usepackage{bbm}

\theoremstyle{plain}
\usepackage{amsthm}
\makeatletter
\newcommand{\newreptheorem}[2]{\newtheorem*{rep@#1}{\rep@title}\newenvironment{rep#1}[1]{\def\rep@title{#2 \ref*{##1}}\begin{rep@#1}}{\end{rep@#1}}}
\makeatother

\newtheorem{theorem}{Theorem}
\newtheorem*{theorem-non}{Theorem}
\newtheorem*{non-lemma}{Lemma}
\newtheorem{lemma}[theorem]{Lemma}
\newreptheorem{lemma}{Lemma}

\newtheorem{claim}[theorem]{Claim}

\newtheorem{conjecture}[theorem]{Conjecture}
\newtheorem{problem}[theorem]{Problem}

\theoremstyle{definition}

\DeclareMathOperator{\Sub}{Sub}

\DeclareMathOperator{\RowS}{RowSpace}

\DeclareMathOperator{\cok}{cok}

\DeclareMathOperator{\supp}{supp}
\DeclareMathOperator{\Sur}{Sur}
\DeclareMathOperator{\Hom}{Hom}
\DeclareMathOperator{\Aut}{Aut}

\DeclareMathOperator{\Sg}{Sub}

\begin{document}
\title{A phase transition for the cokernels of random band matrices over the p-adic integers}
\author{Andr\'as M\'esz\'aros}
\date{}
\affil{HUN-REN Alfr\'ed R\'enyi Institute of Mathematics}
\maketitle
\begin{abstract}
Let $\mathbf{B}_n$ be an $n\times n$ Haar-uniform band matrix over $\mathbb{Z}_p$ with band width $w_n$. We prove that $\cok(\mathbf{B}_n)$ has Cohen-Lenstra limiting distribution if and only if \[\lim_{n\to\infty} \left(w_n-\log_p(n)\right)=+\infty.\]

\end{abstract}

\section{Introduction}

Let $p$ be a prime. Given two positive integers $n$ and $w$, let $\mathcal{B}_{n,w}$ be the set of $n\times n$ band matrices over the $p$-adic integers $\mathbb{Z}_p$ with band width $w$, that is, let 
\[\mathcal{B}_{n,w}=\{B\in M_n(\mathbb{Z}_p)\,:\,B(i,j)=0\text{ for all }i,j\text{ such that }|i-j|>w\}.\]

Let $w_n$ be a sequence of positive integers, and let $\mathbf{B}_n$ be a Haar uniform element of $\mathcal{B}_{n,w_n}$. The cokernel of $\mathbf{B}_n$ is the random abelian $p$-group
\[\cok(\mathbf{B}_n)=\mathbb{Z}_p^n/\RowS(\mathbf{B}_n),\]
where $\RowS(\mathbf{B}_n)$ is the $\mathbb{Z}_p$-submodule of $\mathbb{Z}_p^n$ generated by the rows of $\mathbf{B}_n$. 

Our main theorem states that $\cok(\mathbf{B}_n)$ exhibits a phase transition. If $w_n$ grows fast enough, then  $\cok(\mathbf{B}_n)$ has Cohen-Lenstra limiting distribution. If $w_n$ does not grow fast enough, then $\cok(\mathbf{B}_n)$ does not have Cohen-Lenstra limiting distribution. We determine precisely where this phase transition happens.

\begin{theorem}\hfill\label{thmmain}
\begin{enumerate}[(a)]
\item If $\lim_{n\to\infty} \left(w_n-\log_p(n)\right)=+\infty$, then $\cok(\mathbf{B}_n)$ has Cohen-Lenstra limiting distribution, that is, for all finite abelian $p$-groups $G$, we have
\begin{equation}\label{CohenLenstra}\lim_{n\to\infty} \mathbb{P}(\cok(\mathbf{B}_n)\cong G)=\frac{1}{|\Aut(G)|}\prod_{j=1}^{\infty}\left(1-p^{-j}\right).\end{equation}
\item If $w_n-\log_p(n)$ does not converge to $+\infty$, then $\cok(\mathbf{B}_n)$ does not have Cohen-Lenstra limiting distribution.
\end{enumerate}
\end{theorem}

The distribution on the right hand side of \eqref{CohenLenstra} is called the Cohen-Lenstra distribution. It first appeared in a conjecture about the distribution of class groups of quadratic number fields~\cite{cohen2006heuristics}. We prove the first part of Theorem~\ref{thmmain} by using the moment method of Wood~\cite{wood2017distribution}. By now the moment method is considered the standard method of establishing the Cohen-Lenstra limiting distribution of cokernels of random matrices. See \cite{wood2019random,nguyen2022random,recent1,recent2,recent3,recent4,recent5,recent6,recent7,recent8,recent9,recent10,meszaros2020distribution,meszaros2023cohen} for some recent examples. See also \cite{sawin2022moment,van2024symmetric,wood2022probability} for further results on the moment problem. With the choice of $w_n=n$, Theorem~\ref{thmmain} gives back the classical result of Friedman and Washington~\cite{friedman1987distribution} on the cokernels of Haar uniform matrices over $\mathbb{Z}_p$. Theorem~\ref{thmmain} provides us random matrices with only $O(n\log n)$ non-zero entries and Cohen-Lenstra limiting behaviour. Note that the author provided random matrices with $O(n)$ non-zero entries and Cohen-Lenstra limiting behaviour~\cite{meszaros2020distribution,meszaros2023cohen}. Excluding these examples all other results mentioned above are about denser random matrices. 

The second part of Theorem~\ref{thmmain} is proved by showing that for small band widths, the kernel of the mod $p$ reduction of $\mathbf{B}_n$ contains several localized vectors with sufficiently large probability, see Theorem~\ref{thmloc}. This phenomenon can be viewed as an analogue of the localization/delocalization phase transition of Gaussian band matrices as we explain in Section~\ref{secGaussian}. 

\subsection{Band matrices over $\mathbb{F}_p$}

Let $\overline{\mathbf{B}}_n\in M_n(\mathbb{F}_p)$ be the mod $p$ reduction of $\mathbf{B}_n$. In other words, $\overline{\mathbf{B}}_n$ is a uniform random element of the set 
\[\{B\in M_n(\mathbb{F}_p)\,:\,B(i,j)=0\text{ for all }i,j\text{ such that }|i-j|>w_n\}.\]

The first part of Theorem~\ref{thmmain} has the following corollary.

\begin{theorem}\label{thmrank}
Assume that $\lim_{n\to\infty} w_n-\log_p(n)=+\infty$. Then for all nonnegative integer $k$, we have
\[\lim_{n\to\infty} \mathbb{P}(\dim \ker \overline{\mathbf{B}}_n=k)=p^{-k^2} \prod_{j=1}^{k} \left(1-p^{-j}\right)^{-2} \prod_{j=1}^{\infty}\left(1-p^{-j}\right).\]
\end{theorem}

Next we consider the case when $w_n-\log_p(n)$ does not converge to $+\infty$.

Given two positive integers $n$ and $\lC $, we set $n_0=\left\lfloor\frac{n}\lC \right\rfloor$, and for $i=1,2,\dots,\lC$, we define
\[V_i=\left\{v\in \mathbb{F}_p^n\,:\,\emptyset\neq \supp v\subset \left[(i-1)n_0+1\,,\,i n_0\right]\right\}.\]

We say that $\overline{\mathbf{B}}_n$ has an $\lC $-localized kernel if the following event occurs:
\[\text{For all }i=1,2,\dots,\lC, \text{ there is a }v_i\in V_i\text{ such that }\overline{\mathbf{B}}_n v_i=0.\]

Observe that for $i\neq j$ and $v_i\in V_i,v_j\in V_j$, the vectors $v_i$ and $v_j$ do not interact in the sense that the events  $\overline{\mathbf{B}}_n v_i=0$ and $\overline{\mathbf{B}}_n v_j=0$ depend on disjoint subsets of the entries of~$\overline{\mathbf{B}}_n$. In particular, these two events are independent.  

Note that if $\overline{\mathbf{B}}_n$ has an $\lC $-localized kernel, then $\dim \ker \overline{\mathbf{B}}_n\ge \lC $.

\begin{theorem}\label{thmloc}
Assume that $ w_n-\log_p(n)$ does not converge to $+\infty$. Then there is a $\kappa>0$ with the property that for all positive integers $\lC $, we have infinitely many $n$ such that
\[\mathbb{P}(\,\overline{\mathbf{B}}_n\text{ has an $\lC $-localized kernel}\,)\ge \left(\frac{\kappa}\lC \right)^{4\lC }.\]
Consequently, for all positive integer $\lC $, we have infinitely many $n$ such that
\begin{equation}\label{heavytail}\mathbb{P}(\dim\ker \overline{\mathbf{B}}_n\ge \lC )\ge \left(\frac{\kappa}\lC \right)^{4\lC }.
\end{equation}
\end{theorem}

Note that \eqref{heavytail} excludes the possibility of a Cohen-Lenstra limiting distribution for $\cok(\mathbf{B}_n)$. Thus, the second part of Theorem~\ref{thmmain} follows.

When $w_n$ grows fast enough, we do not see localized vectors in $\ker \overline{\mathbf{B}}_n$. In fact, $\ker \overline{\mathbf{B}}_n$ only consists of typical vectors in the following sense.

\begin{theorem}\label{thmdeloc}
Assume that $\lim_{n\to\infty} w_n-\log_p(n)=+\infty$. Let $\mathcal{E}_n$ be a deterministic subset of  $\mathbb{F}_p^n\setminus\{0\}$ such that $\lim_{n\to\infty} \frac{|\mathcal{E}_n|}{p^n}=0$. Then
\[\lim_{n\to\infty}\mathbb{P}(\ker \overline{\mathbf{B}}_n\cap \mathcal{E}_n\neq\emptyset)=0.\]
\end{theorem}

We prove two further phase transitions.

\begin{theorem}\label{thmtight}
The sequence of random variables $\dim\ker \overline{\mathbf{B}}_n$ is tight if and only if \[\liminf_{n\to\infty} \left(w_n-\log_p(n)\right)>-\infty.\]
\end{theorem}

 It is also very natural to normalize  $\dim\ker \overline{\mathbf{B}}_n$ by $n$. The next theorem gives a phase transition for this normalized dimension.

\begin{theorem}\label{thmnormalized}
 The sequence of random variables \[\frac{\dim\ker \overline{\mathbf{B}}_n}n\]
 converge to $0$ in probability if and only if $\lim_{n\to\infty} w_n=\infty.$
 \end{theorem}
\subsection{Comparison with Gaussian band matrices }\label{secGaussian}
An $n\times n$ Gaussian band matrix with band width $w$ is a random matrix of the form,
\[X_{n,w}=\frac{1}{\sqrt{w}}\begin{pmatrix}
d_1&a_{1,2}&a_{1,3}&\cdots&\cdots& a_{1,w}&0&\cdots&\cdots&0\\
\overline{a_{1,2}}&d_2&a_{2,3}&a_{2,4}&\cdots&\cdots&a_{2,w+1}&0&\cdots& 0\\
\overline{a_{1,3}}&\overline{a_{2,3}}&\ddots&\ddots& & & & & & \vdots\\
\vdots&\overline{a_{2,4}}&\ddots& \ddots& \ddots& & & & & \vdots\\
\vdots&\vdots& &\ddots&\ddots&\ddots& & &  &\vdots\\
\overline{a_{1,w}}&\vdots& & &\ddots&\ddots& \ddots& &  &\vdots\\
0&\overline{a_{2,w+1}}& & & &\ddots &\ddots &\ddots &&\vdots\\
\vdots&0& & & & &\ddots &\ddots &\ddots &\vdots\\
\vdots&\vdots& & & &&  &\ddots &\ddots &\vdots\\
0&0& \cdots&\cdots &\cdots &\cdots &\cdots &\cdots &\cdots&\\

\end{pmatrix},\]
where $(d_i)_{1\le i\le n}$ and $(a_{ij})_{1\le i\le n,\,\,  i<j\le\min(i+w-1,n)}$ are independent standard real and complex Gaussian variables, respectively.

It is conjectured that band matrices exhibit a phase transition \cite{bourgade2018random}:
\begin{itemize}
    \item If $w\ll \sqrt{n}$, then the eigenvectors of $X_{n,w}$ are localized, and the local eigenvalue statistics converge to a Poisson process. 
    \item If $w\gg \sqrt{n}$, then the eigenvectors of $X_{n,w}$ are delocalized, and the local eigenvalue statistics converge to the $\text{Sine}_2$ process. 
\end{itemize}

The localization of eigenvectors was proved for $w\ll n^{1/8}$ in \cite{schenker2009eigenvector}. Later the exponent $1/8$ was improved to ${1/7}$ and then to ${1/4}$ \cite{peled2019wegner,cipolloni2024dynamical,chen2022random}. For fixed $w$, Poisson eigenvalue statistics were proved in \cite{brodie2022density}, see also  \cite{hislop2022local} for further progress towards Poisson statistics.

Improving earlier results \cite{band1,band2,band3,band4,band5}, the delocalization of eigenvectors and the convergence to the $\text{Sine}_2$ process were proved in \cite{bb1,bb2,bb3} for $w\gg n^{3/4}$.

Recent results \cite{van2023local,van2023p,recent5,van2021limits,neretin2013hua,bufetov2017ergodic,assiotis2022infinite} seem to indicate that there is some analogy between the spectrum of classical random matrix ensembles and the cokernels of random matrices over the $p$-adic integers. Theorem~\ref{thmmain} gives a $p$-adic analogue of the Poisson/$\text{Sine}_2$ phase transition. Theorems~\ref{thmloc}~and~\ref{thmdeloc} give a $p$-adic analogue of the localization/delocalization phase transition.

\medskip

\textbf{The structure of the paper:} In Section~\ref{sec2}, we discuss the moment problem for abelian $p$-groups, and explain that in order to prove the first part of Theorem~\ref{thmmain}, it is enough to show that all the surjective moments of $\cok(\mathbf{B}_n)$ converge to $1$. We also show that the first part of Theorem~\ref{thmmain} implies Theorem~\ref{thmrank}, and that Theorem~\ref{thmloc} implies the second part of Theorem~\ref{thmmain}. In Section~\ref{secmomentcalc}, we show that assuming that $\lim_{n\to\infty} w_n-\log_p(n)=+\infty$ all the surjective moments of $\cok(\mathbf{B}_n)$ converge to $1$, which completes the proof of the first part of Theorem~\ref{thmmain}. We finish Section~\ref{secmomentcalc} by the proof of Theorem~\ref{thmdeloc}. Theorems~\ref{thmloc},~\ref{thmtight} and~\ref{thmnormalized} are proved in Sections~\ref{secloc},~\ref{sectight} and~\ref{secnorm}, respectively. In Section~\ref{secopen}, we discuss several open problems. 

\medskip

\textbf{Acknowledgement:} 
The author was supported by the KKP 139502 project.

\section{From convergence of moments to converge of distributions}\label{sec2}

Given two finite abelian groups $G$ and $H$, let $\Hom(G,H)$ and $\Sur(G,H)$ be the set of homomorphisms and surjective homomorphisms from $G$ to $H$, respectively. The next theorem was first proved in \cite{ellenberg2016homological}, and then in greater generality in \cite{wood2017distribution}.

\begin{theorem}\label{thmWood}
    Let $\Gamma_n$ be a sequence of random finite abelian $p$-groups such that for any deterministic finite abelian $p$-group $G$, we have
    \[\lim_{n\to\infty} \mathbb{E}|\Sur(\Gamma_n,G)|=1.\]
    Then for any finite abelian $p$-group $G$, we have
    \[\lim_{n\to\infty} \mathbb{P}(\Gamma_n\cong G)=\frac{1}{|\Aut(G)|}\prod_{j=1}^{\infty}\left(1-p^{-j}\right).\]

\end{theorem}

The expectation $\mathbb{E}|\Sur(\Gamma_n,G)|$ is called the surjective $G$-moment of $\Gamma_n$. Our next lemma gives an expression for $\mathbb{E}|\Sur(\Gamma_n,G)|$, when $\Gamma_n=\cok(M_n)$ for some $n\times n$ random matrix $M_n$ over~$\mathbb{Z}_p$. Let $G$ be a finite abelian $p$-group. Note that $G$ is a $\mathbb{Z}_p$-module. So for all $\mathbf{g}\in G^n$, we can define $M_n \mathbf{g}\in G^n$ the usual way. The map $\mathbf{g}\mapsto M_n\mathbf{g}$ is a $\mathbb{Z}_p$-module homomorphism from $G^n$ to $G^n$. For the proof of the next lemma, see for example \cite[Proposition 5.2.]{meszaros2020distribution}. 

\begin{lemma}\label{lemma6}
We have
\begin{align*}\mathbb{E}|\Hom(\cok(M_n),G)|&=\mathbb{E}|\{\mathbf{g}\in G^n\,:\,M_n\mathbf{g}=0\}|,\text{ and}\\\mathbb{E}|\Sur(\cok(M_n),G)|&=\mathbb{E}|\{\mathbf{g}\in G^n\,:\,M_n\mathbf{g}=0\text{ and the components of $\mathbf{g}$ generate }G\}|.
\end{align*}
\end{lemma}

For two integers $c\le d$, by $[c,d]$ we mean the set $\{c,c+1,\dots,d\}$, and $[c,c-1]$ is defined to be the empty set.  

For an interval $[c,d]$ and a boundary condition 
\[\mathbf{b}:[c-w,c-1]\cup [d+1,d+w]\to G,\] let
\[G^{[c,d]}_\mathbf{b}=\{\mathbf{g}:[c-w,d+w]\to G\,:\,\mathbf{g}(k)=\mathbf{b}(k)\text{ for all }k\in [c-w,c-1]\cup [d+1,d+w]\}.\]

Let $\Sg(G)$ be the set of subgroups of $G$. For any $\mathbf{g}\in G^{[c,d]}_\mathbf{b}$, we define $\mathbf{H}_\mathbf{g}:[c,d]\to\Sg(G)$ by setting
\[\mathbf{H}_\mathbf{g}(k)=\left\langle \mathbf{g}(j)\,:\, k-w\le j\le k+w\right\rangle\]
for all $k\in [c,d]$. 

Let ${0}_n$ be the all zero map defined on $[1-w_n,0]\cup[n+1,n+w_n]$. For $\mathbf{g}\in G^{[1,n]}_{{0}_n}$, let
\[\langle \mathbf{g}\rangle=\langle \mathbf{g}(k)\,:\,k\in [1,n]\rangle.\]

\begin{lemma}\label{lemmamoment}
We have
\begin{align*}\mathbb{E}|\Hom(\cok(\mathbf{B}_n),G)|&=\sum_{\mathbf{g}\in G^{[1,n]}_{{0}_n}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\text{ and }\\\mathbb{E}|\Sur(\cok(\mathbf{B}_n),G)|&=\sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\langle\mathbf{g}\rangle=G}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}.
    \end{align*}

\end{lemma}
\begin{proof}
Let $\tilde{\mathbf{g}}\in G^{[1,n]}$. By appending $\tilde{\mathbf{g}}$ with zeros, we obtain a $\mathbf{g}\in G^{[1,n]}_{{0}_n}$.

Consider the homomorphism $\varphi:\mathcal{B}_{n,w_n}\to G^n$ defined by $\varphi(B)=B\tilde{\mathbf{g}}$. The image of $
\varphi$ is $H=\bigoplus_{k=1}^n \mathbf{H}_\mathbf{g}(k)$. Since $\varphi$ is a homomorphism, it pushes forward the Haar-measure to the Haar-measure. Thus, $\mathbf{B}_n\tilde{\mathbf{g}}$ is a uniform random element of $H$, so 
\begin{equation}\label{Pker}\mathbb{P}(\mathbf{B}_n\tilde{\mathbf{g}}=0)=\prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}.\end{equation}
Combining this with Lemma~\ref{lemma6}, the statement follows.
\end{proof}

In Section~\ref{secmomentcalc}, we will prove the following theorem.
\begin{theorem}\label{thmmoment}
Assume that $\lim_{n\to\infty} w_n-\log_p(n)=+\infty$. Then
\[\lim_{n\to\infty} \sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\langle\mathbf{g}\rangle=G}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}=1.\]
\end{theorem}

The first part of Theorem~\ref{thmmain} follows by combining Theorem~\ref{thmWood}, Lemma~\ref{lemmamoment} and Theorem~\ref{thmmoment}. Theorem~\ref{thmrank} follows from Theorem~\ref{thmmain} and the following lemma.

\begin{lemma}\label{CLrank}
Let $M_n$ be a sequence of random matrices over $\mathbb{Z}_p$, and let $\overline{M}_n$ be the mod $p$ reduction of $M_n$. Assume that for all finite abelian $p$-group $G$, we have
\[\lim_{n\to\infty} \mathbb{P}(\cok(M_n)\cong G)=\frac{1}{|\Aut(G)|}\prod_{j=1}^{\infty}\left(1-p^{-j}\right).\]

Then for any nonnegative integer $k$, we have
\[\lim_{n\to\infty} \mathbb{P}(\dim \ker \overline{M}_n=k)=p^{-k^2} \prod_{j=1}^{k} \left(1-p^{-j}\right)^{-2} \prod_{j=1}^{\infty}\left(1-p^{-j}\right).\]

\end{lemma}
\begin{proof}
See \cite[Theorem 6.3]{cohen2006heuristics}.
\end{proof}

It follows easily from Lemma~\ref{CLrank} that \eqref{heavytail} excludes the possibility of a Cohen-Lenstra limiting distribution for $\cok(\mathbf{B}_n)$, see \cite[Section 5]{meszaros20242} for more details. Therefore, Theorem~\ref{thmloc} implies the second part of Theorem~\ref{thmmain}.

\section{Calculation of moments}\label{secmomentcalc}

Let $t=|\Sg(G)|$, and let $G_1,G_2,\dots,G_t$ be a list of all the subgroups of $G$ such that \break $|G_1|\ge |G_2|\ge\cdots \ge |G_t|$. Clearly, $G_1=G$ and $G_t=\{0\}$.

For $\mathbf{g}\in G^{[c,d]}_\mathbf{b}$, we define
\[\ell(\mathbf{g})=\max\{i\,:\,\mathbf{H}_{\mathbf{g}}(k)=G_i\text{ for some }k\in [c,d]\}.\]

We also define
\begin{align*}G^{[c,d]}_\mathbf{b}[i]&=\{\mathbf{g}\in G^{[c,d]}_\mathbf{b}\,:\,\ell(\mathbf{g})=i\},\text{ and }\\
G^{[c,d]}_\mathbf{b}[\le i]&=\cup_{j=1}^i G^{[c,d]}_\mathbf{b}[j].\end{align*}

Let $n, w$ be positive integers. We set $q=tn p^{-w}$. For $1\le i\le t$, we define
\[\alpha_i=\alpha_i(q)=\begin{cases}
1&\text{for $i=1$,}\\
(q\beta_i+2t)^2\exp(q^2\beta_i)&\text{otherwise,}
\end{cases}\]
where $\beta_i=\sum_{j=1}^{i-1} \alpha_j$.

\begin{lemma}\label{lemmaalpha}
For all $[c,d]$ such that $d-c+1\le n$, $\mathbf{b}:[c-w,c-1]\cup [d+1,d+w]\to G$ and $1\le i\le t$, we have
\[\sum_{\mathbf{g}\in G^{[c,d]}_\mathbf{b}[i]} \prod_{k=c}^d |\mathbf{H}_\mathbf{g}(k)|^{-1}\le \alpha_i.\]

\end{lemma}
\begin{proof}
We prove by induction on $i$. First consider the case $i=1$. By definition, we have $\mathbf{H}_g(k)=G$ for all $k\in [c,d]$ and $\mathbf{g}\in G^{[c,d]}_\mathbf{b}[1]$. Thus,
\[\sum_{\mathbf{g}\in G^{[c,d]}_\mathbf{b}[1]} \prod_{k=c}^d |\mathbf{H}_\mathbf{g}(k)|^{-1}=\frac{|\mathbf{g}\in G^{[c,d]}_\mathbf{b}[1]|}{|G|^{d-c+1}}\le 1.\]

Now assume that $i>1$. Given a $\mathbf{g}\in G^{[c,d]}_\mathbf{b}[i]$, let
\[\mathbb{W}(\mathbf{g})=\{k\in [c,d]\,:\,\mathbf{H}_\mathbf{g}(k)=G_i\}\]
and let \[\mathcal{W}_i=\{\mathbb{W}(\mathbf{g})\,:\, \mathbf{g}\in G^{[c,d]}_\mathbf{b}[i]\}.\]

For a $W\in \mathcal{W}_i$, let 
\[\mathbb{W}^{-1}(W)=\{\mathbf{g}\in G^{[c,d]}_\mathbf{b}[i]\,:\,\mathbb{W}(\mathbf{g})=W\}.\]
\begin{claim}\label{cl13}
Let $x,y\in [c,d]$ such that $x\le y\le x+2w+1$, and let $\mathbf{g}\in G^{[c,d]}_\mathbf{b}[i]$. If $x,y\in \mathbb{W}(\mathbf{g})$, then $[x,y]\subset \mathbb{W}(\mathbf{g})$.

\end{claim}
\begin{proof}
Let $z\in [x,y]$, then $[z-w,z+w]\subset [x-w,x+w]\cup [y-w,y+w]$. Thus, $\mathbf{H}_\mathbf{g}(z)$ is a subgroup of the subgroup generated by $\mathbf{H}_\mathbf{g}(x)$ and $\mathbf{H}_\mathbf{g}(y)$. By our assumptions, $\mathbf{H}_\mathbf{g}(x)=\mathbf{H}_\mathbf{g}(y)=G_i$. Thus, $\mathbf{H}_\mathbf{g}(z)$ is a subgroup of $G_i$. Since $\ell(\mathbf{g})=i$, we see that $|\mathbf{H}_\mathbf{g}(z)|\ge |G_i|$. Therefore, it follows that $\mathbf{H}_\mathbf{g}(z)=G_i$.  
\end{proof}

Claim~\ref{cl13} gives that any $W\in \mathcal{W}_i$ can be uniquely written as 
\[W=\cup_{j=1}^h [c_j,d_j]\]
for $h=h(W)$ subintervals of $[c,d]$ such that  \begin{equation}\label{eqsep}c_{j+1}>d_j+2w+1\text{ for all }1\le j<h.\end{equation}

Let \[L_j=[c_j-w,c_j-1]\cap [c,d]\quad\text{ and }\quad R_j=[d_j+1,d_j+w]\cap [c,d].\] Note that we have $L_j=[c_j-w,c_j-1]$ for $1<j\le h$ and $L_1=[\max(c,c_1-w),c_1-1]$. Similarly, we have $R_j=[d_j+1,d_j+w]$ for $1\le j< h$ and $R_h=[d_h+1,\min(d_h+w,d)]$. Note that these intervals are all pairwise disjoint by \eqref{eqsep}. 

For a $\mathbf{g}\in \mathbb{W}^{-1}(W)$, let
\[H_j^L(\mathbf{g})=\left\langle \mathbf{g}(k)\,:\,k\in L_j\right\rangle\quad\text{ and }\quad H_j^R(\mathbf{g})=\left\langle \mathbf{g}(k)\,:\,k\in R_j\right\rangle.\]
It is easy to see that $H_j^L(\mathbf{g})$ is a subgroup of $\mathbf{H}_\mathbf{g}(c_j)=G_i$.  Similarly, $H_j^R(\mathbf{g})$ is a subgroup of~$G_i$.
\begin{claim}\label{Claim12} Let $W\in \mathcal{W}_i$  
and $\mathbf{g}\in \mathbf{W}^{-1}(W)$. For all $k\in L_j$, we have $|\mathbf{H}_\mathbf{g}(k)|\ge p|H_j^L(\mathbf{g})|$. 

Similarly, for all $k\in R_j$, we have $|\mathbf{H}_\mathbf{g}(k)|\ge p|H_j^R(\mathbf{g})|$.
\end{claim}
\begin{proof}
Let $k\in L_j$. Since $\ell(\mathbf{g})=i$, we get $|\mathbf{H}_\mathbf{g}(k)|\ge |G_i|$. Since $k\notin W$, we have $\mathbf{H}_\mathbf{g}(k)\neq G_i$. Thus, $\mathbf{H}_\mathbf{g}(k)\cap G_i\neq \mathbf{H}_\mathbf{g}(k)$. Observe that $H_j^L(\mathbf{g})$ is a subgroup of $\mathbf{H}_\mathbf{g}(k)\cap G_i=\mathbf{H}_\mathbf{g}(k)\cap \mathbf{H}_\mathbf{g}(c_j)$, because $L_j$ is contained in both $[k-w,k+w]$  and $[c_j-w,c_j+w]$.  
 
 Therefore, $p|H_j^L(\mathbf{g})|\le p |\mathbf{H}_\mathbf{g}(k)\cap G_i|\le |\mathbf{H}_\mathbf{g}(k)|$. The second statement follows similarly.
\end{proof}

For $W\in\mathcal{W}_i$, let $h=h(W)$. Let $F^L_1,F^R_1,\dots, F_h^L,F_h^R$ be subgroups of $G_i$. We define
\[\mathcal{A}(W,F^L_1,F^R_1,\dots, F_h^L,F_h^R)=\{\mathbf{g}\in \mathbb{W}^{-1}(W)\,:\,H_j^L(\mathbf{g})=F^L_j, H_j^R(\mathbf{g})=F_j^R\text{ for all }1\le j\le h\}. \]

Let $B=B(W)=W\cup [c-w,c-1]\cup [d+1,d+w]\cup \cup_{j=1}^h (L_j\cup R_j)$.

Let 
\[\mathcal{B}(W,F^L_1,F^R_1,\dots, F_h^L,F_h^R)=\{\mathbf{g}\restriction B\,:\,\mathbf{g}\in \mathcal{A}(W,F^L_1,F^R_1,\dots, F_h^L,F_h^R)\}\subset G^B. \]

Clearly,
\begin{equation}\label{BWSize}|\mathcal{B}(W,F^L_1,F^R_1,\dots, F_h^L,F_h^R)|\le |G_i|^{|W|}\prod_{j=1}^h \left(|F_j^L|^{|L_j|}\cdot |F_j^R|^{|R_j|}\right).\end{equation}

For $\mathbf{f}\in \mathcal{B}(W,F^L_1,F^R_1,\dots, F_h^L,F_h^R)$, let
\[\mathcal{A}(W,\mathbf{f})=\{\mathbf{g}\in \mathcal{A}(W,F^L_1,F^R_1,\dots, F_h^L,F_h^R)\,:\,\mathbf{g}\restriction B=\mathbf{f}\}.\]

\begin{claim}\label{Claim13}
    Assume that $c_1> c+w$ and $d_h< d-w$. Then for any $\mathbf{f}\in \mathcal{B}(W,F^L_1,F^R_1,\dots, F_h^L,F_h^R)$, we have
    \[\sum_{\mathbf{g}\in \mathcal{A}(W,\mathbf{f})} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le |G_i|^{-|W|} \beta_i^{h+1}p^{-2hw}\prod_{j=1}^h (|F_j^L||F_j^R|)^{-w}.\]
\end{claim}
\begin{proof}
Let $[\overline{c}_0,\overline{d}_0]=[c,c_1-w-1]$, $[\overline{c}_h,\overline{d}_h]=[d_h+w+1,d]$, and for $0<j<h$, let $[\overline{c}_j,\overline{d}_j]=[d_j+w+1,c_{j+1}-w-1]$. 

For any $\mathbf{g}\in \mathcal{A}(W,\mathbf{f})$, we have
\begin{align*}\prod_{k=c}^d &|\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\\&=\left( \prod_{k\in W} |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\right)\left(\prod_{j=1}^h \prod_{k\in L_i} |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\right)\left(\prod_{j=1}^h \prod_{k\in R_i} |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\right)\left(\prod_{j=0}^h \prod_{k\in [\overline{c}_j,\overline{d}_i]} |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\right)\\&\le |G_i|^{-|W|} 
    \left(\prod_{j=1}^h p^{-w}|F_j^L|^{-w} \right)\left(\prod_{j=1}^h p^{-w}|F_j^R|^{-w} \right)\left(\prod_{j=0}^h \prod_{k\in [\overline{c}_j,\overline{d}_i]} |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\right)\end{align*}
where we used Claim~\ref{Claim12}.

 Let $\mathbf{b}_j=\mathbf{f}\restriction ([\overline{c}_j-w,\overline{c}_j-1]\cup [\overline{d}_j+1,\overline{d}_j+w])$. For any $\mathbf{g}\in \mathcal{A}(W,\mathbf{f})$ and $0\le j\le h$, we have $\mathbf{g}\restriction [\overline{c}_j-w,\overline{d}_j+w]\in G^{[\overline{c}_j,\overline{d}_j]}_{\mathbf{b}_j}[\le i-1]$. Therefore,
 \begin{align*}
\sum_{\mathbf{g}\in \mathcal{A}(W,\mathbf{f})} &\prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\\&\le  |G_i|^{-|W|} 
    \left(\prod_{j=1}^h p^{-w}|F_j^L|^{-w} \right)\left(\prod_{j=1}^h p^{-w}|F_j^R|^{-w} \right)\sum_{\mathbf{g}\in \mathcal{A}(W,\mathbf{f})}\prod_{j=0}^h \prod_{k\in [\overline{c}_j,\overline{d}_i]} |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\\&\le |G_i|^{-|W|} p^{-2wh} \left(\prod_{j=1}^h (|F_j^L||F_j^R|)^{-w}\right)\left(\prod_{j=0}^h\sum_{\mathbf{g}\in G^{[\overline{c}_j,\overline{d}_j]}_{\mathbf{b}_j}[\le i-1] }\prod_{k=\overline{c}_j}^{\overline{d}_j} |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\right) \\&\le |G_i|^{-|W|} p^{-2wh} \left(\prod_{j=1}^h (|F_j^L||F_j^R|)^{-w}\right)  \left(\sum_{j=1}^{i-1}\alpha_j\right)^{h+1},
\end{align*}
where in the last step we used the induction hypothesis. Recalling that $\beta_i=\sum_{j=1}^{i-1}\alpha_j$, the statement follows.
\end{proof}

\begin{claim}\label{Claim14}
If $c_1>c+w$ and $d_h<d-w$, then
\[\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le p^{-2hw} t^{2h}\beta_i^{h+1}.\]  
\end{claim}
\begin{proof}

Combining Claim~\ref{Claim13} with \eqref{BWSize}, we obtain that for any choice of $F^L_1,F^R_1,\dots, F_h^L,F_h^R$, we have

\[\sum_{\mathbf{g}\in \mathcal{A}(W,F^L_1,F^R_1,\dots, F_h^L,F_h^R)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le p^{-2hw} \beta_i^{h+1}.\]
    Summing these over all possible choices of $F^L_1,F^R_1,\dots, F_h^L,F_h^R$, we get the statement.
\end{proof}

With a similar proof, one gets:
\begin{claim}
If $c_1\le c+w$ and $d_h<d-w$, then
\[\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le p^{-(2h-1)w-(c_1-c)} t^{2h}\beta_i^{h}.\]
If $c_1>c+w$ and $d_h\ge d-w$, then
\[\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le p^{-(2h-1)w-(d-d_h)} t^{2h}\beta_i^{h}.\]
If $c_1\le c+w$ and $d_h\ge d-w$, then
\[\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le p^{-(2h-2)w-(d-d_h)-(c_1-c)} t^{2h}\beta_i^{h-1}.\]
\end{claim}

Assuming that $h(W)=h$, we have at most ${{d-c+1}\choose{h}}\le \frac{n^h}{h!}$ choices for $c_1,c_2,\dots,c_h$. The same is true for $d_1,d_2,\dots,d_h$. Thus,
\begin{equation}\label{Wibound}|\{W\in \mathcal{W}_i\,:\, h(W)=h\}|\le \left(\frac{n^h}{h!}\right)^2\le \frac{n^{2h}}{(h-1)!}.\end{equation}
Similarly, for any $x,y\in [c,d]$, we have
\begin{align*}|\{W\in \mathcal{W}_i\,:\, h(W)=h, c_1(W)=x\}|&\le \frac{n^{2h-1}}{(h-1)!},\\
|\{W\in \mathcal{W}_i\,:\, h(W)=h, d_h(W)=y\}|&\le \frac{n^{2h-1}}{(h-1)!},\\
|\{W\in \mathcal{W}_i\,:\, h(W)=h, c_1(W)=x,d_h(W)=y\}|&\le \frac{n^{2h-2}}{(h-1)!}.
\end{align*}

We have
\begin{align}\label{sump1}\sum_{\substack{W\in\mathcal{W}_i,\\ c_1(W)>c+w,\\d_{h(W)}(W)<d-w}}&\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\\&=\sum_{h=1}^\infty \sum_{\substack{W\in\mathcal{W}_i,\\h(W)=h,\\c_1(W)>c+w,\\d_{h}(W)<d-w}}\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}_{\mathbf{g}}(k)|^{-1}\nonumber\\&\le \sum_{h=1}^\infty \frac{n^{2h}}{(h-1)!}
      p^{-2hw} t^{2h}\beta_i^{h+1}\nonumber\\&=q^2\beta_i^2 \sum_{h=1}^\infty \frac{\left(q^2\beta_i\right)^{h-1}}{(h-1)!}\nonumber\\&=q^2\beta_i^2\exp(q^2\beta_i),\nonumber
\end{align}
where at the first inequality, we used Claim~\ref{Claim14} and \eqref{Wibound}.

We also have
\begin{align}\label{sump2}\sum_{\substack{W\in\mathcal{W}_i,\\ c_1(W)\le c+w,\\d_{h(W)}(W)<d-w}}&\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\\&=\sum_{h=1}^\infty \sum_{u=0}^w\sum_{\substack{W\in\mathcal{W}_i\\ h(W)=h,\\c_1(W)=c+u,\\d_{h}(W)<d-w}}\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}_{\mathbf{g}}(k)|^{-1}\nonumber\\&\le \sum_{u=0}^{w}\sum_{h=1}^\infty \frac{n^{2h-1}}{(h-1)!}
      p^{-(2h-1)w-u} t^{2h}\beta_i^{h}\nonumber\\&=\left(\sum_{u=0}^w p^{-u}\right)tq\beta_i \sum_{h=1}^\infty \frac{\left(q^2\beta_i\right)^{h-1}}{(h-1)!}\nonumber\\&\le 2tq\beta_i\exp(q^2\beta_i).\nonumber
\end{align}

Similarly,
\begin{equation}\label{sump3}
    \sum_{\substack{W\in\mathcal{W}_i,\\ c_1(W)> c+w,\\d_{h(W)}(W)\ge d-w}}\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le 2tq\beta_i\exp(q^2\beta_i).
\end{equation}

Finally,
\begin{align}\label{sump4}\sum_{\substack{W\in\mathcal{W}_i,\\ c_1(W)\le c+w,\\d_{h(W)}(W)\ge d-w}}&\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\\&=\sum_{h=1}^\infty \sum_{u=0}^w \sum_{v=0}^w\sum_{\substack{W\in\mathcal{W}_i,\\ h(W)=h,\\c_1(W)=c+u,d_{h}(W)=d-v}}\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}_{\mathbf{g}}(k)|^{-1}\nonumber\\&\le \sum_{u=0}^{w}\sum_{v=0}^{w}\sum_{h=1}^\infty \frac{n^{2h-2}}{(h-1)!}
      p^{-(2h-2)w-u-v} t^{2h}\beta_i^{h-1}\nonumber\\&=\left(\sum_{u=0}^w p^{-u}\right)^2t^2\sum_{h=1}^\infty \frac{\left(q^2\beta_i\right)^{h-1}}{(h-1)!}\nonumber\\&=4t^2\exp(q^2\beta_i).\nonumber
\end{align}

For later use, we record that almost the same argument gives that
\begin{equation}\label{sump4b}\sum_{\substack{W\in\mathcal{W}_i\\h(W)>1\\ c_1(W)\le c+w,\\d_{h(W)}(W)\ge d-w}}\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le 4t^2\left(\exp(q^2\beta_i)-1\right).\end{equation}
Combining \eqref{sump1},\eqref{sump2},\eqref{sump3} and \eqref{sump4}, we obtain Lemma~\ref{lemmaalpha}.
\end{proof}

Let
\[\mathcal{W}^*_i=\mathcal{W}_i\cap \{[c_1,d_1]\,:\,c\le c_1\le c+w, d-w\le d_1\le d\}.\]

Combining  \eqref{sump1},\eqref{sump2},\eqref{sump3} and \eqref{sump4b}, we see that
\begin{equation}\label{sumpc}\sum_{W\in\mathcal{W}_i\setminus \mathcal{W}^*_i}\sum_{\mathbf{g}\in \mathbb{W}^{-1}(W)} \prod_{k=c}^d |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le q\beta_i(4t+q\beta_i)\exp(q^2\beta_i)+4t^2\left(\exp(q^2\beta_i)-1\right).\end{equation}


\begin{lemma}\label{lemmamomentupper}
Assume that $\lim_{n\to\infty}w_n-\log_p(n)=+\infty$,
then 
\[\limsup_{n\to\infty} \sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\langle\mathbf{g}\rangle=G}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le 1.\]

\end{lemma}
\begin{proof}
Let $q_n=tnp^{-w_n}$. Clearly, $\lim q_n=0$. Since $\beta_i(q)$ is monotone increasing in $q$, there is a $C$ such that $\beta_i(q_n)<C$ for all $n$ and $2\le i\le t$. It follows from \eqref{sumpc} that
\begin{align*}\sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\W(\mathbf{g})\notin \mathcal{W}_{\ell(\mathbf{g})}^*}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}&\le \sum_{i=2}^t \left( q_n\beta_i(q_n)(4t+q_n\beta_i(q_n))\exp(q_n^2\beta_i(q_n))+4t^2\left(\exp(q_n^2\beta_i(q_n))-1\right)\right)\\&\le (t-1)\left( q_nC(4t+q_nC)\exp(q_n^2C)+4t^2\left(\exp(q_n^2C)-1\right)\right). \end{align*}
Here the right hand side converges to $0$. Therefore, 
\[\limsup_{n\to\infty} \sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\langle\mathbf{g}\rangle=G}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le \limsup_{n\to\infty} \sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\langle\mathbf{g}\rangle=G,\mathbb{W}(\mathbf{g})\in \mathcal{W}^*_{\ell(\mathbf{g})}}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}.\]

Observe that for any $\mathbf{g}\in G^{[1,n]}_{{0}_n}$, we have $\mathbf{H}_{\mathbf{g}}(1)\subset \mathbf{H}_{\mathbf{g}}(2)\subset \cdots \subset \mathbf{H}_{\mathbf{g}}(w_n+1)$ and $\mathbf{H}_{\mathbf{g}}(n)\subset \mathbf{H}_{\mathbf{g}}(n-1)\subset \cdots \subset \mathbf{H}_{\mathbf{g}}(n-w_n)$. Thus, if $\mathbb{W}(\mathbf{g})\in \mathcal{W}^*_{\ell(\mathbf{g})}$, then it follows that $\mathbb{W}(\mathbf{g})=[1,n]$. Thus, $\mathbf{H}_\mathbf{g}(k)=G_{\ell(\mathbf{g})}$ for all $k\in [1,n]$. If we further require that $\langle \mathbf{g}\rangle=G$, then it follows that $\mathbf{H}_\mathbf{g}(k)=G$ for all $k\in [1,n]$. Therefore,

\begin{equation}\label{onlytypical}
    \lim_{n\to\infty} \left(\sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\langle\mathbf{g}\rangle=G}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\quad \mathlarger{-}\sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\mathbf{H}_\mathbf{g}(k)=G\text{ for all }k\in[1,n]}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\right)=0.
\end{equation}

Thus,
\[\limsup_{n\to\infty} \sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\langle\mathbf{g}\rangle=G}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le \limsup_{n\to\infty} \sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\mathbf{H}_\mathbf{g}(k)=G\text{ for all }k\in[1,n]}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le |G|^n|G|^{-n}=1.
\]

\end{proof}

Observe that \begin{align}\label{sq0}\limsup_{n\to\infty} \sum_{\substack{\mathbf{g}\in G^{[1,n]}_{{0}_n}\\\langle\mathbf{g}\rangle=G}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}&\ge \limsup_{n\to\infty}(|G|^n-\sum_{G\neq G'\in \Sg(G)} |G'|^n)|G|^{-n}\\&\ge \limsup_{n\to\infty} (1-|\Sg(G)|p^{-n})=1.\nonumber\end{align}

Combining this with Lemma~\ref{lemmamomentupper}, we obtain Theorem~\ref{thmmoment}.

We finish this section by proving Theorem~\ref{thmdeloc}.

Given a $\tilde{\mathbf{g}}\in G^{[1,n]}$, let  $\mathbf{g}\in G^{[1,n]}_{{0}_n}$  be obtained by appending $\tilde{\mathbf{g}}$ with zeros.

Let \[\mathcal{E}_n'=\{\tilde{\mathbf{g}}\in \mathcal{E}_m\,:\,\mathbf{H}_\mathbf{g}(k)=\mathbb{F}_p\text{ for all }k\in[1,n]\}.\]

Note that since $\mathbb{F}_p$ is simple group, $\langle \mathbf{g}\rangle=\mathbb{F}_p$ for all nonzero $\mathbf{g}$. Thus, under the assumption that $\lim_{n\to\infty} w_n-\log_p(n)\to\infty$, it follows from \eqref{onlytypical} that
\[\limsup_{n\to\infty} \sum_{\tilde{\mathbf{g}}\in \mathcal{E}_n} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\le \limsup_{n\to\infty} \sum_{\tilde{\mathbf{g}}\in \mathcal{E}_n'}\prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}.\]

Combining this with the union bound, \eqref{Pker} and the definition of $\mathcal{E}_n'$, we see that
\begin{align*}
\limsup_{n\to\infty}\mathbb{P}(\ker \overline{\mathbf{B}}_n\cap \mathcal{E}_n\neq\emptyset)&\le\limsup_{n\to\infty}\sum_{\tilde{\mathbf{g}}\in \mathcal{E}_n} \mathbb{P}(\mathbf{B}_n\tilde{\mathbf{g}}=0)\\&=\limsup_{n\to\infty} \sum_{\tilde{\mathbf{g}}\in \mathcal{E}_n} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\\&\le \limsup_{n\to\infty} \sum_{\tilde{\mathbf{g}}\in \mathcal{E}_n'}\prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\\&=\limsup_{n\to\infty}|\mathcal{E}_n'|\cdot p^{-n}=0.
\end{align*}
Thus, Theorem~\ref{thmdeloc} follows.
\section{Localization}\label{secloc}

If $w_n-\log_p(n)$ does not converge to $+\infty$, then there is a $1>\varepsilon>0$ such that $np^{-w_n}\ge\varepsilon$ for infinitely many $n$. Thus, Theorem~\ref{thmloc} follows from the lemma below.

\begin{lemma}\label{loclemma}
Let $\varepsilon>0$ and $\lC\ge \varepsilon $ be an integer. Choose $w\ge 1$ and $n> 2\lC$ such that $np^{-w}\ge \varepsilon$. Let $\mathbf{B}_n$ be a Haar-uniform element of $\mathcal{B}_{n,w}$. Set $n_0= \left\lfloor\frac{n}\lC \right\rfloor$. 

Let
\[V_i=\left\{v\in \mathbb{F}_p^n\,:\,\emptyset\neq \supp v\subset \left[(i-1)n_0+1\,,\,i n_0\right]\right\}.\]

Let us consider the random variables
\[X_i=|\{v\in V_i\,:\, \mathbf{B}_nv=0 \}|.\]

Then these random variables are independent and
\[\mathbb{P}(X_i\ge 1)\ge \left(C\frac{\varepsilon}\lC \right)^{4}\text{ for all }i=1,2,\dots,\lC\text{ and }\]
\[\mathbb{P}(X_i\ge 1\text{ for all }i=1,2,\dots,\lC )\ge \left(C\frac{\varepsilon}\lC \right)^{4\lC }\]
for some positive constant $C$ only depending on $p$.

\end{lemma}
\begin{proof}
Let us set $n_1=\min\left(n_0,p^{w}\right)$.
Using the assumptions $n>2L$ and $np^{-w}\ge \varepsilon$, we have 
\begin{equation}\label{qbound}\frac{n_1}{p^w}= \min\left(\left\lfloor\frac{n}{\lC}\right\rfloor\frac{1}{p^w},1\right)\ge \min\left(\frac{n}{2\lC p^w},1\right)\ge \min\left(\frac{\varepsilon}{2\lC },1\right)=\frac{\varepsilon}{2\lC }, \end{equation}
where in the last step we used $\lC\ge\varepsilon$.

Let $(i-1)n_0+1\le c<d\le (i-1) n_0+n_1$, and let
\[U=\{v\in \mathbb{F}_p^n\,:\,\{c,d\}\subset \supp v\subset [c,d]\}.\]

Note that
\begin{equation}\label{Ubound}|U|=(p-1)^2p^{d-c-1}\ge p^{d-c-1}.\end{equation}
For all $v\in U$, the distribution of $\mathbf{B}_nv$ is uniform on some subgroup of 
\[U'=\{v\in \mathbb{F}_p^n\,:\, \supp v\subset [c-w,d+w].\}\]

Thus,
\[\mathbb{P}(\mathbf{B}_nv=0)\ge p^{-(d-c+1+2w)}.\]
Combining this with \eqref{Ubound}, we obtain
\begin{equation}\label{EUb}\mathbb{E}|\{v\in U\,:\,\mathbf{B}_nv=0\}|\ge p^{-2w-2}.\end{equation}

We define
\[V_i'=\left\{v\in \mathbb{F}_p^n\,:\,\emptyset\neq \supp v\subset \left[(i-1)n_0+1\,,\,(i-1) n_0+n_1\right]\right\},\]
and the random variables
\[Y_i=|\{v\in V_i'\,:\, \mathbf{B}_nv=0 \}|.\]

Clearly, $V_i'\subset V_i$, and $Y_i\le X_i$.
Then using \eqref{EUb}, we get
\[\mathbb{E}Y_i\ge \sum_{(i-1)n_0+1\le c<d\le (i-1)n_0+n_1} p^{-2w-2}={{n_1}\choose{2}}p^{-2w-2}\ge \frac{n_1^2 p^{-2w}}{4p^2},\]
using the fact that $n_1>1$ so $n_1-1\ge \frac{n_1}2$.

Let $\mathbf{B}'$ be the submatrix of $\mathbf{B}_n$ determined by the columns and rows with index in the interval $[a,z]=[(i-1)n_0+1-w\,,\,(i-1) n_0+n_1+w]\cap [1,n]$. Note $\mathbf{B}'$ has the same distribution as an $n_2\times n_2$ Haar-uniform band matrix with band width $w$, where
\[n_2=\min((i-1) n_0+n_1+w,n)-\max((i-1) n_0+1-w,1)+1\le n_1+2w\le 3p^w.\]

Then
\[\mathbb{E} Y_i^2=\mathbb{E}|\{(v,v')\in (V_i')^2\,:\,\mathbf{B}_nv=0,\mathbf{B}_nv'=0 \}|\le \mathbf{E}|\{u\in (\mathbb{F}_p\times 
\mathbb{F}_p)^{n_2}\,:\,\mathbf{B}'u=0 \}|\le \gamma\]
for some $\gamma$. The existence of $\gamma$ follows from Lemma~\ref{lemmaalpha} and the fact that $n_2 p^{-w}\le 3$. (Note that $\alpha_i$ is monotone increasing in $q$.) The second to last inequality follows from the observation that $\mathbf{B}_nv=0,\mathbf{B}_nv'=0$ if and only if $\mathbf{B}'u=0$, where $u\in (\mathbb{F}_p\times \mathbb{F}_p)^{n_2}$ such that \[u_j=\left(v_{a+j-1},v'_{a+j-1}\right).\]

From the Paley-Zygmund inequality,
\begin{equation}\label{paley}\mathbb{P}(X_i\ge 1)\ge\mathbb{P}(Y_i\ge 1)\ge \frac{(\mathbb{E} Y_i)^2}{\mathbb{E} Y_i^2}\ge \frac{n_1^4 p^{-4w}}{16p^4\gamma}\ge \left(C\frac{\varepsilon}\lC \right)^4,\end{equation}
where $C=\frac{1}{4p\sqrt[4]{\gamma}}$. At the last inequality we used \eqref{qbound}.

Note that for all $v\in V_i$ the event $\mathbf{B}_n v=0$ on the depend on the columns of $\mathbf{B}_n$  indexed by $[(i-1)n_0+1,in_0]$. Thus, $X_1,\dots,X_\lC $ are independent. Combining this with  \eqref{paley} , we obtain
\[\mathbb{P}(X_i\ge 1\text{ for all }i=1,2,\dots,\lC )\ge  \left(C\frac{\varepsilon}\lC \right)^{4\lC }.\]
\end{proof}

\section{Tightness}\label{sectight}
In this section, we prove Theorem~\ref{thmtight}.

First assume that $\liminf_{n\to\infty} w_n-\log_p(n)>-\infty$. Let $q_n=2np^{-w_n}$. Then there is a $q$ such that $q_n\le q$ for all $n$. Combining Lemma~\ref{lemmamoment} and Lemma~\ref{lemmaalpha}, we see that
\begin{align*}
\mathbb{E}|\ker \overline{\mathbf{B}}_n|&=\mathbb{E}|\Hom(\cok(\mathbf{B}_n),\mathbb{F}_p)|\nonumber\\&=\sum_{\mathbf{g}\in G^{[1,n]}_{{0}_n}} \prod_{k=1}^n |\mathbf{H}
    _{\mathbf{g}}(k)|^{-1}\nonumber\\&\le 1+(q_n+4)^2\exp\left(q_n^2\right)\\&\le 1+(q+4)^2\exp(q^2).\nonumber
\end{align*}

Thus, $\sup \mathbb{E}|\ker \overline{\mathbf{B}}_n|<\infty$, which by Markov's inequality implies that the sequence $|\ker \overline{\mathbf{B}}_n|$ is tight. Therefore, the sequence $\dim \ker \overline{\mathbf{B}}_n=\log_p |\ker \overline{\mathbf{B}}_n|$ is also tight.

Now, assume that $\liminf_{n\to\infty} w_n-\log_p(n)=-\infty$.

Let $\lC$ be a positive integer and set $\varepsilon=\lC$. There are infinitely many $n$ such that $n p^{-w_n}\ge \varepsilon$ and $n> 2\lC$. For any such $n$, Lemma~\ref{loclemma} can be applied to give that $\mathbb{P}(X_i\ge 1)\ge C^4$. Let $I_i$ be the indicator of the event that $X_i\ge 1$. By Lemma~\ref{loclemma}, $I_1,I_2,\dots,I_\lC$ are independent and
\[\mathbb{E}\sum_{i=1}^{\lC} I_i\ge \lC C^4.\]
It follows from Hoeffding inequality that
\[\mathbb{P}\left(\dim \ker \overline{\mathbf{B}}_n\le \frac{\lC C^4}2\right)\le \mathbb{P}\left(\sum_{i=1}^{\lC} I_i\le \frac{\lC C^4}2\right)\le \exp\left(-\frac{C^8\lC}{2}\right),\]
which clearly makes it impossible that the sequence $\dim \ker \overline{\mathbf{B}}_n$ is tight. 

 \section{Normalized dimension}\label{secnorm}
 In this section, we prove Theorem~\ref{thmnormalized}.

First assume that $\lim_{n\to\infty} w_n=\infty.$ Given $n$, let us write $[1,n]$ as the disjoint union of $h=\lceil\frac{n}{w_n}\rceil$ intervals $I_1,\dots,I_h$ such that $|I_j|=n_j\le w_n.$ Let $\mathbf{M}_j$ be the $n_j\times n_j$ submatrix of $\overline{\mathbf{B}}_n$ determined by the rows and columns indexed by $I_j$. Since $n_j\le w_n$, we see that $\mathbf{M}_j$ is a uniform random $n_j\times n_j$ matrix over $\mathbb{F}_p$. Finally, let $\mathbf{C}_n$ be obtained from $\overline{\mathbf{B}}_n$ by setting to $0$ all the entries in the region $\cup_{j=1}^h I_j\times I_j$. Note that even if we condition on $\mathbf{C}_n$, $\mathbf{M}_j$ is still a uniform random $n_j\times n_j$ matrix over $\mathbb{F}_p$ and $\mathbf{M}_1,\dots,\mathbf{M}_h$ are conditionally independent. Given a vector $v\in \mathbb{F}_p^n$, let $v_{I_j}\in \mathbb{F}_p^{n_j}$ be its restriction to $I_j$.

Note that $\mathbf{B}_n v=0$ if and only if $\mathbf{M}_j v_{I_j}=-(\mathbf{C}_nv)_{I_j}$ for all $1\le j\le h$. Using the conditional independence of $\mathbf{M}_1,\dots,\mathbf{M}_h$, we obtain
\[\mathbb{P}(\mathbf{B}_n v=0\,|\,\mathbf{C}_n)=\prod_{j=1}^h \mathbb{P}(\mathbf{M}_j v_{I_j}=-(\mathbf{C}_nv)_{I_j}\,|\,\mathbf{C}_n). \]

If $v_{I_j}\neq 0$, then $\mathbf{M}_j v_{I_j}$ even conditioned on $\mathbf{C}_n$ is a uniform random element of $\mathbb{F}_p^{n_j}$. Thus, in this case
\[\mathbb{P}(\mathbf{M}_j v_{I_j}=-(\mathbf{C}_nv)_{I_j}\,|\,\mathbf{C}_n)=p^{-n_j}.\]
Therefore,
\[\mathbb{P}(\mathbf{B}_n v=0)\le \prod_{\substack{1\le j\le h\\v_{I_j}\neq 0}} p^{-n_j}.\]
So
\[\mathbb{E}|\ker \mathbf{B}_n|=\sum_{v\in\mathbb{F}_p^n} \mathbb{P}(\mathbf{B}_n v=0)\le \prod_{j=1}^h(1+(p^{n_j}-1)p^{-n_j})\le 2^h\le 2^{n/w_n+1}.\]

Combining this with Markov's inequality,  for all $\varepsilon>0$, we obtain that
\begin{multline*}\limsup_{n\to\infty}\mathbb{P}(\dim\ker \overline{\mathbf{B}}_n\ge \varepsilon n)=\limsup_{n\to\infty}\mathbb{P}(|\ker \overline{\mathbf{B}}_n|\ge p^{\varepsilon n}) 
\\\le\limsup_{n\to\infty}\frac{\mathbb{E}|\ker \overline{\mathbf{B}}_n|}{p^{\varepsilon n}}\le\limsup_{n\to\infty}2^{n/w_n+1-\varepsilon n}=0. \end{multline*}

Thus,
 \[\frac{\dim \ker \overline{\mathbf{B}}_n}n\]
 converge to zero in probability.

Now assume that $w_n$ does not converge to infinity. Then there is a finite $w$ such that $w_n=w$ infinitely often. Consider such an $n$. Let $Z_i$ be the indicator of the event that the $i$th column of $\overline{\mathbf{B}}_n$ is the all zero vector. We have $\mathbb{E}{Z}_i\ge p^{-2w-1}$, so  $\mathbb{E}\sum_{i=1}^n{Z}_i\ge p^{-2w-1}n$. Clearly, $Z_1,\dots,Z_n$ are independent, thus by Hoeffding's inequality, we see that
\[\mathbb{P}\left(\dim \ker\overline{\mathbf{B}}_n\le \frac{1}2 p^{-2w-1}n\right)\le \mathbb{P}\left(\sum_{i=1}^{n} Z_i\le \frac{1}2p^{-2w-1}n\right)\le \exp\left(-\frac{1}2 p^{-4w-2}n\right),\]
which shows that \[\frac{\dim \ker \overline{\mathbf{B}}_n}n\]
 can not converge to zero in probability.

\section{Open problems}\label{secopen}

\subsection{Behaviour at criticality}

Our tightness result given in Theorem~\ref{thmtight} points towards the following conjecture.
\begin{conjecture}
  There is a one parameter family of distributions $(\nu_{\alpha})_{\alpha\in\mathbb{R}}$ on the set of finite abelian $p$-groups with the following property. Let $n_1<n_2<\dots$ be a sequence of positive integers. Let $\mathbf{B}_i$ be an $n_i\times n_i$ band matrix over $\mathbb{Z}_p$ with band width $w_i$. Let us assume that $\lim_{i\to\infty} w_i-\log_p(n_i)=\alpha$. Then for a finite abelian $p$-group $G$, we have
  \[\lim_{i\to\infty} \mathbb{P}(\cok(\mathbf{B}_i)\cong G)=\nu_\alpha(G).\]
\end{conjecture}
   Provided that this conjecture is true, $\nu_\alpha$ must have a tail which is heavier than the tail of the Cohen-Lenstra distribution as we can see from Theorem~\ref{thmloc}. Note that if the $2$-torsion of the homology of determinatal hypertrees has a limiting distribution, then this limiting distribution also has a tail which is heavier than the tail of the Cohen-Lenstra distribution \cite{meszaros20242}.

\subsection{Symmetric band matrices}

\begin{problem}
    Can we prove an analogue of Theorem~\ref{thmmain} for symmetric band matrices? The Cohen-Lenstra distribution should be replaced with a modified version of the Cohen-Lenstra distribution given in \cite{clancy2015note,clancy2015cohen}.
\end{problem}

\subsection{General entry distributions}

Let $\tau$ be a probability measure on $\mathbb{Z}_p$ which is non-degenerate in the sense that $\tau(i+p\mathbb{Z}_p)<1$ for all $i=0,1,\dots,p-1.$ Consider band matrices where the non-zero entries are i.i.d. with law~$\tau$.

\begin{problem}
    Can we prove an analogue of Theorem~\ref{thmmain} for such a matrices band matrices? Is it true that the phase transition happens at $c(\tau)\log_p(n)$, where $c(\tau)$ is some function of $\tau$.
\end{problem}

\subsection{General prescription matrices}

By a prescription matrix $P$, we mean an $n\times n$ matrix, where each entry is either $0$ or $1$. The set of $P$-band matrices is defined as
\[\mathcal{B}_{P}=\{B\in M_n(\mathbb{Z}_p)\,:\,B(i,j)=0\text{ for all }i,j\text{ such that }P(i,j)=0\}.\]

Let $\mathbf{B}_P$ a Haar uniform element of $\mathcal{B}_{P}$. 

We say a sequence $P_n$ of prescription matrices is weakly Cohen-Lenstra, if the distribution of $\cok(\mathbf{B}_{P_n})$ converges to the Cohen-Lenstra distribution. We say a sequence $P_n$ of prescription matrices is strongly Cohen-Lenstra, if $\lim_{n\to\infty} \mathbb{E}|\Sur(\cok(\mathbf{B}_{P_n}),G)|=1$ for all finite abelian $p$-group $G$. It follows from Theorem~\ref{thmWood} that if $P_n$ is strongly Cohen-Lenstra, then it is weakly Cohen-Lenstra.

\begin{problem}
    It is true that $(P_n)$ is weakly Cohen-Lenstra if and only if strongly Cohen-Lenstra?
\end{problem}

\begin{problem}\label{problem22}
    Try to characterize the weakly/strongly Cohen-Lenstra sequences. 
\end{problem}

Note that Theorem~\ref{thmmain} provides such a characterization in the special case when $P_n$ is an $n\times n$ matrix such that
\[P_n(i,j)=\begin{cases}1&\text{if $|i-j|\le w$},\\0&\text{otherwise}.\end{cases}\]

A natural and interesting special case Problem~\ref{problem22} is the following:
\begin{problem}
Find the critical band width for $d$-dimensional band matrices.
\end{problem}

Note that the set of strongly Cohen-Lenstra sequences of prescription matrices is upwardly closed. More precisely, we have the following lemma.

\begin{lemma}
Let $(P_n)$ and $(Q_n)$ be two sequences of prescription matrices such that for all $n$, $P_n$ and $Q_n$ has the same dimensions, and $P_n\le Q_n$ entrywise.  If $P_n$ is strongly Cohen-Lenstra, then $Q_n$ is also strongly Cohen-Lenstra.
\end{lemma}
\begin{proof}
Let $P$ be an $n\times n$ prescription matrix, and
let $G$ be a finite abelian $p$-group. For $\mathbf{g}\in G^n$, let us define $\mathbf{H}_{\mathbf{g},P}:[1,n]\to \Sg(G)$ by
\[\mathbf{H}_{\mathbf{g},P}(k)=\langle \mathbf{g}(j)\,:\, P(k,j)=1\rangle.\]

Then $\mathbf{B}_P\mathbf{g}$ is a uniform random element of $\bigoplus_{k=1}^n \mathbf{H}_{\mathbf{g},P}(k)$. Thus, as in the proof Lemma~\ref{lemmamoment}, we see that
\[\mathbb{E}|\Sur(\cok(\mathbf{B}_P,G))|=\sum_{\substack{\mathbf{g}\in G^n\\\langle\mathbf{g}\rangle=G}}\prod_{k=1}^n |\mathbf{H}_{\mathbf{g},P}(k)|^{-1}, \]
where as before $\langle\mathbf{g}\rangle$ is the subgroup generated by the components of $\mathbf{g}.$

If $P\le Q$ entrywise, then $\mathbf{H}_{\mathbf{g},P}(k)$ is a subgroup of $\mathbf{H}_{\mathbf{g},Q}(k)$ for any $\mathbf{g}\in G^n$ and $1\le k\le n$. Thus, it follows that
\[\mathbb{E}|\Sur(\cok(\mathbf{B}_P,G))|\ge \mathbb{E}|\Sur(\cok(\mathbf{B}_Q,G))|\ge 1-|\Sg(G)|p^{-n},\]
where the last inequality follows the same way as \eqref{sq0}. The lemma follows from the squeeze theorem.
\end{proof}

The next problem asks whether Theorem~\ref{thmmain} provides us the asymptotically sparsest weakly (or strongly) Cohen-Lenstra sequence of prescription matrices.
\begin{problem}
  Let $(P_n)$ be a weakly/strongly Cohen-Lenstra sequence of prescription matrices such that $P_n$ is an $n\times n$ matrix. Is it true that 
  \[\liminf_{n\to\infty}\frac{1}{n\log_p(n)}\sum_{i=1}^n\sum_{j=1}^n P_n(i,j)\ge 2?\]
\end{problem} 
\bibliography{references}
\bibliographystyle{plain}

\bigskip

\bigskip

\noindent Andr\'as M\'esz\'aros, \\
HUN-REN Alfr\'ed R\'enyi Institute of Mathematics, \\Budapest, Hungary,\\ {\tt meszaros@renyi.hu}
\end{document}